ItzJasminefools:

Since tweedle dee and tweedle dumb wanted to get annoying i actually do have a question and I will post it in the chat

TETSXPREME:

? where?

ItzJasminefools:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @TETSXPREME ? where? \(\color{#0cbb34}{\text{End of Quote}}\) GIVE ME TIME ITS LOADING

ItzJasminefools:

2x + 3(4x−6)+5

BabydollBlake:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @TETSXPREME ? where? \(\color{#0cbb34}{\text{End of Quote}}\) Right

TETSXPREME:

is this pemdas?

BabydollBlake:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @TETSXPREME is this pemdas? \(\color{#0cbb34}{\text{End of Quote}}\) More of like terms

ItzJasminefools:

No its the one where you have to find the common number

ItzJasminefools:

like 2x and 4x they have varables

TETSXPREME:

ooh

TETSXPREME:

so your supposed to find out what x means in both numbers?

ItzJasminefools:

6x right?

BabydollBlake:

2x+3(4x−6)+5 Distribute =2x+(3)(4x)+(3)(−6)+5 =2x+12x+−18+5 Combine Like Terms =2x+12x+−18+5 =(2x+12x)+(−18+5) =14x+−13

BabydollBlake:

Tell me if you understand that

TETSXPREME:

w\[2x + 3 (4x - 6) + 5 \]

BabydollBlake:

If not ill explain or @TETSXPREME

TETSXPREME:

you have to simplfy it

ItzJasminefools:

Yes I do

BabydollBlake:

Mkay then.

TETSXPREME:

do you know how to ?

ItzJasminefools:

5(x+4)−2x+5 Is 3-10 right cause x+2 is 3 and 5+5 is 10

TETSXPREME:

no

BabydollBlake:

No, you cant add a variable and a number

ItzJasminefools:

Oh so it's 2x+20

TETSXPREME:

i dont think thats right either

BabydollBlake:

Is that your answer??

BabydollBlake:

Cause no it is not right

ItzJasminefools:

oh well then is it OHH I KNOW 5*5=25 its 3+25

BabydollBlake:

good!

TETSXPREME:

yes

BabydollBlake:

But your variable

ItzJasminefools:

sorry 3x+25

BabydollBlake:

any other questions?

ItzJasminefools:

3(4x−2)+2(x+3) Uhm 2x+20?

TETSXPREME:

welp ig im not needed here

BabydollBlake:

No ma'am

BabydollBlake:

thats not correct

ItzJasminefools:

WAIT NO 14x

BabydollBlake:

Hint, one side is gonna zero out, meaning it's only gonna be one number and a variable

BabydollBlake:

yes!

ItzJasminefools:

Yes!! Okay I am going to close this now Thank you btw

BabydollBlake:

No problem!